X²+Y²+Z²-2X+4Y-4Z+k=0 is an equation of sphere, then k could be ........

Grade:12

1 Answers

Vikas TU
6 years ago
By rearranging the equation of sphere
(X-1)2 -1 + (Y+2)2 -4 +(Z-2)2 -4 +k=0
(X-1)2 + (Y+2)2 +(Z-2)2 -9 +k=0
Hence -9 + k = r2    Where r is the radius of the sphere.
Hence r =(k-9)^1/2
Hence k-9 should be positive.
k is greater than 9